NERSCPowering Scientific Discovery Since 1974

March 1

Attending:  Eric and Jay from NERSC; users Andrei, Jeff P., Jeff A., Doug O.

Utilization:  Cluster continues to be filled to capacity.

Outages:  There was an interacxtive reboot; eliza5 maintenance; work on pdsfdtn1 tomorrow; eliza12/13 to go away Friday

Procurements:  New hardware expanded eliza5 by 100%.

Project Accounts:  Getting restarted - Jay to talk to security.

Some LDAP problem were occurring during this meeting...

ATLAS grid:  seems to be working with new entries in the grid-mapfile

STAR/ALICE grid:  Some problems w the ALICE SE;  setting up test w/vo box to send jobs to other gatekeepers

Andrei on cvmfs:  Expect to see more usage, they are no longer using /common; working on getting it on all nodes;  need it on debug and pdsfdtn1 - Jay to look

Jay:  we want to put memory limits on the interactive nodes - probably set to 2GB...

The debug node is dual socket 8 core, 2GB/core so we could consider an upgrade if it starts getting heavier use

Jeff:  fixing cron on ALICE servers - somehow Jeff's access no longer works (?)